# Cash Flow Underwriting with Bank Transaction Data: Advancing MSME Financial Inclusion in Malaysia

### v2

@@ 2. $i$
    Business Demographics: years in business, location, business sector code, number of directors, and director’s minimum age.
@@ 6.2. Quantitative Results
Quantitative results on the validation split of the proposed dataset are shown in Figure 2. The LR model achieves the highest AUROC of 0.782, representing a 30.77% improvement over AB, the lowest-performing model with a score of 0.598. RF follows with 0.655, and GB with 0.633. This ranking can be explained by the fact that LR is well suited to capturing the linear structure of the data, whereas more complex ensemble methods such as RF, GB, and AB are prone to overfitting when applied to relatively small samples, particularly under class imbalance (310 non-event vs. 56 event cases in the training set). In such settings, models that provide well-calibrated probabilities and strong generalization tend to perform better, which is consistent with prior benchmarking studies showing that LR often matches or outperforms tree-based ensembles on small-to-medium credit datasets (Brown and Mues, 2012; Lessmann et al., 2015).
6.3. Ablation Studies
@@
Results in Figure 3 show that bank statement–derived features exhibit stronger predictive power than application form information. Nine of the ten transaction-based features rank above all application features, with the sole exception of customer classification, a pre-assigned label from the lending institution describing the business entity. Customer classification slightly surpasses the recent three-month maximum average balance with a relative difference of 7.92%, while also maintaining a substantial margin of 8.67% over the next highest-ranked application feature. These findings reinforce the superior discriminatory strength of bank statement features in distinguishing default from non-default outcomes.
@@
In addition, we conducted ablation experiments by progressively removing feature groups to study which features matter most for MSMEs credit scoring. We applied 5-fold cross-validation on the training split of the proposed dataset to reduce the risk that results are due to random chance and to increase confidence in model stability. From results illustrated in Figure 4(a), we observe a substantial uplift when incorporating bank transaction data. The AUROC of LR improves from 0.672 (application information only) to 0.821 (bank transaction only), representing an uplift of 0.149 points (approximately 22% relative improvement). When both feature sets are combined, the AUROC of LR further increases to 0.8497, confirming that transactional data provides significant incremental predictive value. The validation results shown in Figure 4(b) are consistent with the cross-validation findings. The LR model achieves the highest AUROC overall, improving from 0.647 (application only) to 0.763 (bank statement only), and further to 0.806 when both feature sets are combined. Similar trends are observed for ensemble models such as RF, GB and AB, which record respective gains of 0.127, 0.145 and 0.181 points with the inclusion of bank transaction data. On both splits, all models that use bank transaction data outperform those based solely on application information. The combined features lead to the best result for all models. In general, these findings strongly support the hypothesis that bank statement transaction data offers significant predictive power in credit risk assessment for MSMEs in Malaysia. While application form information provides some discriminatory power, the inclusion of bank transaction-based features substantially enhances model performance, reinforcing the effectiveness of our proposed workflow.
(a) AUROC performance on the training split with different features.:
(b) AUROC performance on the validation split with different features.:
Figure 4. Comparison of AUROC performance across models and feature sets on the training and validation splits.
7. Deployment and Operational Framework
@@ 7.1. Data Ingestion
The pipeline begins when bank statements are received. Extraction models parse and structure the raw transaction data using Optical Character Recognition (OCR) and layout analysis techniques. This process transforms unstructured documents into a standardized format suitable for downstream processing. The ingestion stage also performs integrity verification and data validation checks to ensure document authenticity.
7.2. Feature Engineering and Feature Store
@@
Following ingestion, structured data are fed into the feature engineering module, which computes transaction-derived and behavioral features. These features are time-stamped, versioned, and stored in a repository. The feature store provides a single source of truth for both training and inference, ensuring the same logic used to generate historical features is consistently applied to new applicants in real time. This design guarantees reproducibility, prevents feature skew, and supports traceability across model iterations.
7.3. Continuous Integration (CI)
@@
The CI pipeline automates model retraining and validation before integration into the production registry. It can be triggered by two key events: (1) on a predefined schedule (e.g., annually), when new labeled data become available from the core banking system, and (2) in response to alerts from the monitoring system indicating model or data drift. When triggered, the pipeline retrieves the latest feature set and ground truth labels, retrains the model, and evaluates it against the deployed model using predefined performance metrics.
7.4. Model Registry and Continuous Deployment (CD)
@@
If the retrained model demonstrates superior performance, it is automatically versioned and stored in the model registry. The CD process then packages the validated model into a containerized microservice and deploys it as a secure REST API endpoint. To ensure reliability and minimize production risk, a canary deployment strategy is applied—initially routing a small fraction of live traffic to the new model for performance validation before full-scale rollout.
7.5. Continuous Monitoring and Feedback Loop
@@
Once deployed, the model operates within a Champion–Challenger framework to enable continuous performance optimization (Kim et al., 2019). The Champion model serves as the current production baseline, while a Challenger model is periodically retrained on the latest data and evaluated in parallel. Their performance is continuously monitored and compared across predefined metrics. When the Challenger consistently outperforms the Champion, the system automatically notifies bank officers to review the results and promote the new model to production. This approach ensures ongoing model improvement, robustness, and operational stability.
7.6. Integrated Credit Scoring Framework
@@
In production, the cash flow underwriting system functions as an integrated decision engine embedded within the lending process. Upon submission of bank statements, AI models automatically extract and structure the raw data using OCR and natural language processing models. The processed data then undergoes fraud detection, cash flow analysis, and network analytics to generate transaction-derived features and credit scores for MSME applicants. The existing bureau-based scorecard, designed for borrowers with established credit histories, operates in parallel with the newly developed cash flow–based model. Each model independently produces a risk rating (Low, Medium, or High). A risk override mechanism ensures conservative decisioning; if either model indicates higher risk, the final classification adopts that rating. This framework provides an evidence-based and data-driven approach to MSME credit assessment using adaptive, explainable AI scoring. It also extends financial access to new-to-credit businesses through verifiable transaction data.
8. Conclusion
@@
In conclusion, this study introduces bank statement transactions as alternative data for MSMEs loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature and practical value of alternative data.
Acknowledgements.
@@
We thank the Cradle Fund and the Microsoft for Startups program for their continuous support. This work was made possible through the contributions of the DocLab and DocSuite teams at AI Lens.

### v3

@@ 2. $i$
    Business Demographics: years in business, business location, business sector code, number of directors, and director’s minimum age.
@@
Quantitative results on the validation split of the proposed dataset are shown in Figure 3. The LR model achieves the highest AUROC of 0.782, representing a 30.77% improvement over AB, the lowest-performing model with a score of 0.598. RF follows with 0.655, and GB with 0.633. This ranking can be explained by the fact that LR is well suited to capturing the linear structure of the data, whereas more complex ensemble methods such as RF, GB, and AB are prone to overfitting when applied to relatively small samples, particularly under class imbalance (310 non-event vs. 56 event cases in the training set). In such settings, models that provide well-calibrated probabilities and strong generalisation tend to perform better, which is consistent with prior benchmarking studies showing that LR often matches or outperforms tree-based ensembles on small-to-medium credit datasets (brown2012experimental; lessmann2015benchmarking).
Ablation Studies
@@
Results in Figure 4(a) show that bank statement–derived features exhibit stronger predictive power than application form information. Nine of the ten transaction-based features rank above all application features, with the sole exception of customer classification, a pre-assigned label from the lending institution describing the business entity. Customer classification slightly surpasses the recent three-month maximum average balance with a relative difference of 7.92%, while also maintaining a substantial margin of 8.67% over the next highest-ranked application feature. These findings reinforce the superior discriminatory strength of bank statement features in distinguishing default from non-default outcomes.
@@
In addition, we conducted ablation experiments by progressively removing feature groups to study which features matter most for MSMEs credit scoring. We applied 5-fold cross-validation on the training split of the proposed dataset to reduce the risk that results are due to random chance and to increase confidence in model stability. From results illustrated in Figure 4(b), we observe a substantial uplift when incorporating bank transaction data. The AUROC of LR improves from 0.672 (application information only) to 0.821 (bank transaction only), representing an uplift of 0.149 points (approximately 22% relative improvement). When both feature sets are combined, the AUROC of LR further increases to 0.8497, confirming that transactional data provides significant incremental predictive value. The validation results are consistent with the cross-validation findings. On both splits, all models that use bank transaction data outperform those based solely on application information. The combined features lead to the best result for all models. In general, these findings strongly support the hypothesis that bank statement transaction data offers significant predictive power in credit risk assessment for MSMEs in Malaysia. While application form information provides some discriminatory power, the inclusion of bank transaction-based features substantially enhances model performance, reinforcing the effectiveness of our proposed workflow.
7 Deployment and Operational Framework
@@
The pipeline begins when bank statements are received. Extraction modules parse and structure the raw transaction data using OCR and layout analysis techniques. This process transforms unstructured documents into a standardised tabular format suitable for downstream processing. The ingestion stage also performs integrity verification and data validation checks to ensure document authenticity.
Feature Engineering and Feature Store
@@
Following ingestion, structured data are fed into the feature engineering module, which computes transaction-derived and behavioural features. These features are time-stamped, versioned, and stored in a centralised repository. The feature store provides a single source of truth for both training and inference, ensuring the same logic used to generate historical features is consistently applied to new applicants in real time. This design guarantees reproducibility, prevents feature skew, and supports traceability across model iterations.
Continuous Integration (CI)
@@
The CI pipeline automates model retraining and validation prior to integration into the production registry. It is triggered by two primary events: (1) a predefined schedule (e.g., annually) when new labelled data become available from the core banking system, and (2) alerts from the monitoring system indicating potential model or data drift. Upon activation, the pipeline retrieves the latest versioned feature set and ground truth labels, retrains the model, and evaluates its performance against the current production model using predefined and consistent evaluation metrics.
Model Registry and Continuous Deployment (CD)
@@
If the retrained model demonstrates superior performance, it is automatically versioned and stored in the model registry. The CD process then packages the validated model into a containerised microservice and deploys it as a secure REST API endpoint. To ensure reliability and minimise production risk, a canary deployment strategy is applied—initially routing a small fraction of live traffic to the new model for performance validation before full-scale rollout.
Continuous Monitoring and Feedback Loop
@@
Once the bank statement-based credit scoring model is deployed, it operates within a Champion–Challenger framework (kim2019champion) to enable continuous performance optimisation. The Champion model serves as the current production baseline, while one or more Challenger models are periodically retrained on the latest data and evaluated in parallel. Model performance is continuously monitored and compared across predefined metrics. When a Challenger consistently outperforms the Champion, the system automatically notifies designated bank officers to review the results and promote the new model to production. This approach ensures sustained model improvement, robustness, and operational stability to maintain optimal performance.
Integrated Credit Scoring Framework
@@
In production, the cash flow underwriting system functions as an integrated decision engine embedded within the lending process. Upon submission of bank statements, AI modules automatically extract and structure the raw data using OCR and natural language processing models. The processed data then undergoes fraud detection, cash flow analysis, and network analytics to generate transaction-derived features and credit scores for MSME applicants. The existing bureau-based scorecard, designed for borrowers with established credit histories, operates in parallel with the newly developed cash flow–based model. Each model independently produces a risk rating, and a risk override mechanism ensures conservative decisioning; if either model indicates higher risk, the final classification adopts that rating. This framework provides a transparent, evidence-based, and data-driven approach to MSME credit assessment using adaptive, explainable AI scoring. It also extends financial access to new-to-credit businesses through verifiable transaction data.
8 Conclusion
@@
This study is the first to introduce bank statement transactions as alternative data for MSME loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature and practical value of alternative data.
9 Acknowledgments
@@
We thank the Cradle Fund and the Microsoft for Startups program for their continuous support. This work was made possible through the collective contributions of the DocLab, DocSuite and MLSuite teams at AI Lens, whose efforts were instrumental in the development of this research.
10 Limitations and Future Work
This work has several limitations that should be addressed in future research. First, the dataset comprises 611 loan applications from a single Malaysian lending institution. While the dataset represents real-world production data collected over a two-year period, the sample size is constrained by the practical realities of MSME lending in emerging markets, where data availability remains limited due to stringent data privacy and regulatory requirements. Future work should validate these findings across multiple institutions and larger datasets as additional MSME lending data become available.
Second, the class imbalance in our dataset (518 non-default vs. 93 default cases) reflects the natural distribution observed in real-world lending portfolios, where the majority of borrowers successfully repay their loans. While this imbalance poses modelling challenges, it accurately reflects real-world credit risk settings. In this work, we addressed this limitation through appropriate model selection, WOE-based feature engineering, and careful evaluation using AUROC, which is robust to class imbalance. It is important to note that class imbalance is not a limitation that can be fully eliminated, but rather an inherent characteristic of credit scoring datasets. As credit assessment models improve and lending decisions become more accurate, successful repayment rates increase, thereby maintaining or even intensifying this natural imbalance. Nevertheless, future studies could explore advanced resampling techniques or cost-sensitive learning methods to further enhance model performance on minority class prediction.
Third, while we describe multiple AI modules for document processing, fraud detection, and transaction analysis within the proposed end-to-end cash flow underwriting workflow, the evaluation focuses on overall credit scoring performance rather than the effectiveness of individual modules. However, detailed module-level assessment is constrained by the use of proprietary methods and sensitive operational data that cannot be publicly disclosed. Future work utilising synthetic or anonymised data could enable more granular module-level analysis and benchmarking. Additionally, future research could explore formalising these modules as fully autonomous agents, enabling the development of a multi-agent credit scoring architecture with clearer inter-agent responsibilities and interactions.
Fourth, the proposed workflow has been deployed in a production environment with one Malaysian lending institution. While this demonstrates real-world applicability, broader validation across different lending institutions, regulatory environments, and MSME segments would strengthen the generalisability of our findings. Further research is needed to systematically integrate the proposed cash flow underwriting workflow with existing credit assessment systems that rely on behavioural and credit bureau data, ensuring coherent and robust risk evaluation. Additionally, longitudinal studies examining model performance stability across economic cycles would provide valuable insights into the temporal robustness and adaptability of bank statement-based features under varying macroeconomic conditions.